OpenStudy (anonymous):

Find the sum of the geometric sequence. three divided by two, three divided by eight, three divided by thirty two, three divided by one hundred and twenty eight, three divided by five hundred and twelve

OpenStudy (aum):

\(\Large S = \frac 32+\frac38+\frac{3}{32}+\frac{3}{128}+...\) \(\Large S = 3(\frac 12+\frac18+\frac{1}{32}+\frac{1}{128}+...)\) What is the common ratio?

OpenStudy (anonymous):

i have no idea :/

OpenStudy (aum):

Divide the second term by the first term.

OpenStudy (anonymous):

.0625

OpenStudy (anonymous):

one divided by two hundred and fifty six one divided by sixteen one thousand twenty three divided by five hundred and twelve 341

OpenStudy (aum):

3/8 divided by 3/2 = 3/8 x 2/3 = 1/4 \(\Large r = \frac 14;~~ a = \frac 32\) \(\Large S = \frac{a}{1-r}\)

OpenStudy (aum):

This is sum of an infinite geometric series, right?

OpenStudy (anonymous):

yeah

OpenStudy (aum):

None of the answer choices you provided matches my answer. Are you sure those are the answer choices for this problem and not for the previous or the next problem?

OpenStudy (aum):

Sum of the infinite geometric series = 2 for this problem.